OpenStudy (anonymous):

x-18=23 solve and check with full workings

OpenStudy (chaise):

x-18=23 +18 +18 x = 41 Assuming x = 41, in the following equation: x-18=23 41-18=23

OpenStudy (anonymous):

You are checking that u added 18 by taking away 18, hadn't you better check that u took away 18 correctly..:-)